OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 8 + . . . + 2n = 2^(n+1) - 2

OpenStudy (shubhamsrg):

p1 mean n=1 same way, pk would mean setting n=k same for k+1 ..

OpenStudy (anonymous):

@shubhamsrg so p1= n which =1 and pk = n? so then pk+1 = n+1?? that's a little confusing..

OpenStudy (shubhamsrg):

hmm,,yep you are confused.. p denotes the statement .. n denotes term number question asks you p1 i.e. statement or expression when n =1 same way, pk would be mean when n= k .. pn = k has no meaning for this context..

OpenStudy (anonymous):

@shubhamsrg do you think you could use them in a simple example?

OpenStudy (shubhamsrg):

leme try let pn be 1 + 2 +3 + 4 .... n p1 would mean when n= 1 ,that is 1st term here, 1st term is 1 ,, so p1 =1 pk would mean kth term nth term is ofcorse = n so kth term will be likewise =k so pk =k for pk+1 ,,just substitute n=k+1 in the nth term..

OpenStudy (anonymous):

@shubhamsrg so for my example, p1= 2 pk = 2n and pk+1 =( 2(n+1) - 2) +1 or 2n+1 ?

OpenStudy (shubhamsrg):

for pk, just substitute k in place of n for pk+1 . just substitute k+1 in place of n p1 is correct though..

OpenStudy (shubhamsrg):

so your pk = k

OpenStudy (shubhamsrg):

tell me pk+1 now..

OpenStudy (anonymous):

@shubhamsrg k+1 ?

OpenStudy (anonymous):

@shubhamsrg can i change all the n's in my problem to k's?

OpenStudy (shubhamsrg):

yes you can,,without the loss of generality..

OpenStudy (anonymous):

@shubhamsrg but what i don't understand is do i have to use the equation on the right? and am i actually plugging in any numbers or anything?

OpenStudy (shubhamsrg):

ohh,,didnt see a RHS there,., :P yes, you are plugging in numbers, natural numbers, in place of n n denotes the term no. like 2 is the first term 4 is the second term. 6 is the 3rd term. and so on so nth term = 2*n thus, you can say, kth term = 2*k k+1 th term = 2(k+1) same way, do it on right hand side,.

OpenStudy (anonymous):

@shubhamsrg so then when you have \[2\times k\] it'll equal \[2\] and when you have \[2k+1 \] it'll equal 3 if \[k=1\]

OpenStudy (sirm3d):

P\(\color {red} 1\): \(2 = 2(\color{red} 1 + 1) - 2\) P\(\color{red}k\): \(2 + 4 + 8 + \cdots + 2\color{red}k = 2(\color{red} k + 1) - 2\)

OpenStudy (anonymous):

@sirm3d which equation do i plug it into?

OpenStudy (sirm3d):

the equation Pn: \(\displaystyle 2 + 4 +8 + \cdots + 2n = 2(n+1) -2\) the problem asks you to write the correct equation when (i) n = 1, (ii) n = k, and (iii) n = k+1

OpenStudy (sirm3d):

P1 means replace \(n\) in Pn by 1 ...

OpenStudy (anonymous):

@sirm3d do i use 2n or 2^(n+1)−2

OpenStudy (sirm3d):

i am just following Pn from the first post, although i think \[\Large Pn: 2 + 4 + 8 + \cdots +2^n=2^{n+1}-2\] is what you intended to write.

OpenStudy (anonymous):

@sirm3d yes , that is what i meant to write. oops.

OpenStudy (sirm3d):

the statement \(Pn\) says that there are several terms to add in the left hand side of the equation, the first three terms of which are 2, 4 and 8, while the right hand side is an expression involving \(n\)

OpenStudy (anonymous):

@sirm3d okay i understand that

OpenStudy (sirm3d):

so we are cleared with P1 and Pk. can you write Pk+1?

OpenStudy (anonymous):

@sirm3d do p1 and pk both equal 2?

OpenStudy (sirm3d):

nope. P1 is specific, n = 1, while Pk is more general, n= k.

OpenStudy (sirm3d):

\[\Large P \color{red}n: 2 + 4 + 8 + \cdots + 2^\color{red}n = 2^\color{red}{n+1} -2 \]\[\Large P \color{red}k: 2 + 4 + 8 + \cdots + 2^\color{red}k = 2^\color{red}{k+1} -2 \]

OpenStudy (sirm3d):

Pk is no different in form from Pn. P1 on the other hand, is more specific. The number of terms to write on the LHS is controlled by (n=1)

OpenStudy (anonymous):

@sirm3d so for p1, you would plug in 1 into the equations?

OpenStudy (sirm3d):

yes.

OpenStudy (anonymous):

@sirm3d so then when you have Pk, how would you do that? :(

OpenStudy (sirm3d):

replace n by k thoughout the equation

OpenStudy (anonymous):

@sirm3d so then it would be \[2k=2^{k=1}-2\]

OpenStudy (sirm3d):

nope. the tricky part is the left hand side of the equation, 2 + 4 + 8 + ... + 2^n

OpenStudy (sirm3d):

to write Pk, copy and paste Pn, replace every instance of n by k.

OpenStudy (sirm3d):

Pn: 2 + 4 + 8 + ... + 2^n = 2^(n+1) - 2 replace n by k in the equation to get you Pk

OpenStudy (sirm3d):

don't forget that the left hand side of Pn has "many" terms, the last of which is \(\Large \displaystyle 2^n\)

OpenStudy (anonymous):

@sirm3d i honestly do not understand this one bit :( i'm sorry :(

OpenStudy (anonymous):

@sirm3d it's not that you're explaining it bad or anything, i just don't understand the concepts

OpenStudy (sirm3d):

let's try by writing P1, P2 P3 and P4 statements

OpenStudy (sirm3d):

can write them out?

OpenStudy (sirm3d):

one at a time.

OpenStudy (anonymous):

@sirm3d p1= 2 because 2x1 =2

OpenStudy (sirm3d):

you should write an equation. the left hand side consists of summands, the right hand side an expression i'll do P1 P1: 2 + 4 + 8 + ... + 2^1 = 2^(1+1) - 2

OpenStudy (sirm3d):

but, wait, the last term in the LHS is 2, so there's only one term on that side P1: 2 = 2^(1 + 1) - 2

OpenStudy (anonymous):

@sirm3d so the equation that i was given has the left hand side as 2n but if you do the math is should be 2^n. hmm p2= \[2^2 = 4 \] = \[2^{(2+1)}-2 = 2^3-2 =8-2=6\] but that is incorrect...

OpenStudy (sirm3d):

here's what you need to remember in this Pn: 2 + 4 + 8 + ... + 2^n = 2^(n+1)-2 there are probably many terms on the left hand side. the first term is 2, and the last term is 2^n. so when n = 1, the first term is 2 and the last term is also 2, there's only one term. now, when n = 2, the first term is 2 and the last term is 2^2=4, so there are two terms.

OpenStudy (sirm3d):

\[P1: 2 = 2^(1+1) -2\]\[P2:2+4=2^(2+1)-2\] it would be a wise guess that in P3, there are three terms to add in the LHS

OpenStudy (anonymous):

@sirm3d but why are there more terms being added to the left side ?

OpenStudy (sirm3d):

that is what Pn asserts: there are terms to add on the LHS, while the RHS gives you a quick way of finding the sum without actually adding them.

OpenStudy (sirm3d):

the LHS of Pn grows as n increases. P1: 2 P2: 2 + 4=(2^2) P3 : 2 + 4 + 8(=2^3) P4: 2 + 4 + 8 + 16(=2^4) P5: 2 + 4 + 8 + 16 + 32(=2^5) P10: 2+4+8+16+32+64+128+256+512+1024(=2^10)

OpenStudy (sirm3d):

the RHS of Pn on the other hand, is just an expression P1: 2^(1+1) -2 P2: 2^(2+1) -2 P3: 2^(3+1) -2 P4: 2^(4+1) -2 P5: 2^(5+1) -2 P10: 2^(10+1)-2

OpenStudy (sirm3d):

\[\Large Pn: \underbrace{2+4+8+\cdots+2^n}_{n\text{ terms}}=\underbrace{2^{n+1}-2}_{\text{one expression}}\]

OpenStudy (sirm3d):

\[\Large P_\color{red}1:\underbrace{?}_{1\text{ term}}=2^{\color{red}1+1}-2\]\[\Large P_\color{red}2:\underbrace{?+?}_{2\text{ terms}}=2^{\color{red}2+1}-2\]\[\Large P_\color{red}3:\underbrace{?+?+?}_{3\text{ terms}}=2^{\color{red}3+1}-2\]\[\Large P_\color{red}4:\underbrace{?+?+?+?}_{4\text{ terms}}=2^{\color{red}4+1}-2\]

OpenStudy (anonymous):

@sirm3d ohhh i understand. omg i get it now. that makes a lot more sense. you add the number from the previous P

OpenStudy (sirm3d):

right.

OpenStudy (anonymous):

@sirm3d so i understand it now. thank you soooooo much for your help!! i honestly appreciate it! have a happy new year :)

OpenStudy (sirm3d):

happy new year to you too. here's a problem if you wish to practice. \[\Large \color{red}{Pn}: \color{blue}{2+8+14+\cdots+(6n-4)=3n^2-n}\]